Help me solve the following:
1. d= 2h.
So, h=?
2. Solve using the elimination method:
3r = 5s = -8
5r + 3s = 66
3. What is the y-intercept?:
y + x = -8

Answer by linh117(31)

1. d=2h
Since h is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation.
2h=d
Divide each term in the equation by 2.
(2h)/(2)=(d)/(2)
Simplify the left-hand side of the equation by canceling the common terms.
h=(d)/(2)
2.
(3r - 5s = -8)(3)
(5r + 3s = 66)(5)
______________________
9r - 15s = -24
25r + 15s = 330
_______________
34r = 306
r=9 plug r in to one of the original equations for s
3(9) - 5s = -8
s=7

3. y-intercept form =
y=mx + b
m=slope
b=y-intercept
y + x = -8
y=-x-8
intercept = -8